Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 02

1. Section 9.4: Problems 3, 5, 7, 9, 11, 13, 15, 19, 23, 25, 28, 34, 35, 39, 40, 41, 44 (Does it converge? Why?).